# Impact of control signal phase noise on qubit fidelity

### v1

@@ Abstract
As qubit decoherence times are increased and readout technologies are improved, nonidealities in the drive signals, such as phase noise, are going to represent a growing limitation to the fidelity achievable at the end of complex control pulse sequencies. Here we study the impact on fidelity of phase noise affecting reference oscillators with the help of numerical simulations, which allow us to directly take into account the interaction between the phase fluctuations in the control signals and the evolution of the qubit state.
Our method is based on the generation of phase noise realizations consistent with a given power spectral density, that are then applied to the pulse carrier in simulations, with Qiskit-Dynamics, of the qubit temporal evolution. By comparing the final state obtained at the end of a noisy pulse sequence with that in the ideal case and averaging over multiple noise realizations, we estimate the resulting degradation in fidelity, and exploiting an approximate analytical representation of a carrier affected by phase fluctuations, we discuss the contributions of the different spectral components of phase noise.
@@ Introduction
Developing large-scale reliable and viable quantum computers is at present one of the main objectives of the scientific community. The physical implementations of the unit of quantum information, the qubit, are characterized by two main parameters, that are fundamental indicators of their performance: the relaxation time $T_{1}$ and the coherence time $T_{2}^{\ast}$ [1, 2, 3, 4]. Both of these parameters must be optimized, and, in particular, decoherence, the process through which freely evolving qubits gradually lose stored information, represents one of the main limitations. Error mitigation techniques, in particular quantum error correction (QEC) algorithms, have been developed, but their successful implementation requires that qubits have coherence times long enough for the QEC routines to be executed [5, 6, 7]. Sources of decoherence are random physical processes in the environment surrounding the quantum processor, but also noise and fluctuations in the signals used to control qubits. Therefore, specifically understanding and reducing the effects of noise on decoherence are fundamental challenges within the overall development of quantum computers.
Here, we focus on the investigation of the action of the noise affecting the pulses used to control qubit evolution, and, in particular, of phase noise, which is the prevalent nonideality (other forms of noise, such as thermal and shot noise, have in general a minor effect in this specific application). We will mainly deal with superconducting qubits, which represent one of the most promising implementations for quantum computers, due to their relatively long coherence times, as well as to the ease of manipulation [8, 9, 10, 11].
@@
Over the years, the need to control quantum states with high precision has led to the development of Quantum Control Theory [12, 13, 14, 15]. An important aspect of this theory is the use of spectral overlap techniques to analyze and reduce the effect of environmental noise on control protocols [16, 17]. This approach considers the spectral characteristics of both the noise, through its power spectral density, and of the control pulses or pulse sequences, through the associated filter functions in the Fourier domain. Based on these assumptions, the theory of Dynamical Error Suppression (DES) was developed, which exploits a sequence of appropriately designed pulses to mitigate the effects of dephasing [18, 19, 20, 21, 22, 23, 24].
@@
A few years ago, the approach based on filter transfer functions was extended to the analysis of the impact, on the coherence and operational fidelity of qubits, of the phase fluctuations (phase noise) associated to the reference oscillators for control pulse generation. In particular, in Ref. [25] H. Ball et al. extended the usage of the filter function formalism to the study of the effect of the phase fluctuations of the reference clock by shifting such fluctuations onto the qubit. Indeed, the choice of the reference frame is arbitrary, and, therefore, one can assume the local oscillator frame as the reference; thus, phase fluctuations of the local oscillator will appear as opposite phase oscillations of the qubit.
@@
Our approach is based on a two-step method: first, we generate a pseudorandom sequence of phase values distributed in agreement with the assumed (or measured) noise Power Spectral Density (PSD); then, such a sequence is used as an input for a numerical simulation of the evolution in time of the qubit state, performed with Qiskit-Dynamics [26]. The final state reached by the qubit can then be compared with the one from an ideal, noiseless evolution. Simulations are repeated for different realizations of the same phase noise process (changing the seed used to generate the pseudorandom sequence) and averaged to obtain an estimate for the fidelity achievable with each considered choice of noise power spectral density.
Results
We use a two-step approach to quantify the impact of control-pulse phase noise on qubit fidelity. This procedure combines the synthetic phase noise generation with the quantum simulation of the evolution in the time domain performed with Qiskit-Dynamics. The method used to generate sequences of baseband phase noise values for the simulations operates in the discrete-time domain and is based on filtering a white Gaussian process through a Finite Impulse Response (FIR) filter, synthesized on the basis of the desired noise PSD and Impulse Response Truncation (IRT). It will be described in detail in the Method section.
Once the baseband phase noise values with specific amplitude and spectral characteristics are obtained, they are used directly as the phase $\phi_{j}$ at time $t_{j}$ of the carrier for the control pulses ($s(t_{j})=\Re\{f(t_{j})\exp[i(2\pi\nu t_{j}+\phi_{j})]\}$, where $f(t_{j})$ is the pulse envelope at $t_{j}$, and $\nu$ is the carrier frequency), in order to simulate their effect on qubit evolution. In all of our simulations we have considered a rectangular envelope for the control pulses.
@@ Discussion
Method
The corresponding finite-length FIR impulse response is then:
$h[n]={}$
In our numerical procedure, we define the amplitude function $A_{d}(\omega)$ as the square root of the desired noise PSD, set $\psi_{0}=0$, and we choose even values for the order $M$ (Type I filter). This numerical procedure, shown in more detail in Fig. 14, allows us to obtain a phase noise process with the desired PSD, just by modeling the shape of the FIR filter frequency response. To accelerate the computation of the convolution in the time domain between the white Gaussian process and the impulse response of the filter, the Fast Fourier Transform (FFT) is leveraged to convert the operation into a multiplication in the frequency domain, performing proper zero-padding to ensure that the vector lengths are powers of 2 [28].

### v2

@@ Abstract
As qubit decoherence times are increased and readout technologies are improved, nonidealities in the drive signals, such as phase noise, are going to represent a crucial limitation to the fidelity achievable at the end of complex control pulse sequences. Although the effect of phase noise of reference oscillators on qubit performance has been studied previously, its interaction with realistic time-dependent control pulses and its contribution to fidelity degradation has not yet been investigated in sufficient detail, and remains a critical challenge. Here we study the impact on fidelity of phase noise affecting reference oscillators with the help of numerical simulations, which allow us to directly take into account the interaction between the phase fluctuations in the control signals and the evolution of the qubit state, thereby achieving a comprehensive understanding of the actual role played by the different spectral components of phase noise. In particular, we perform an analysis of the effect of the individual noise frequency contributions, providing a clear identification of the spectral regions that most critically impact fidelity and establishing their relative weight in the overall fidelity degradation.
Our method is based on the generation of phase noise realizations consistent with a given power spectral density, that are then applied to the pulse carrier in simulations, with Qiskit-Dynamics, of the qubit temporal evolution. By comparing the final state obtained at the end of a noisy pulse sequence with that in the ideal case and averaging over multiple noise realizations, we estimate the resulting degradation in fidelity, and, exploiting an approximate analytical representation of a carrier affected by phase fluctuations, we shed new light on the nature of the different contributions, and provide an intuitive physical picture.
Keywords: phase noise, superconducting qubits, control electronics, fidelity.
@@ Introduction
Developing large-scale reliable and viable quantum computers is at present one of the main objectives of the scientific community. The physical implementations of the unit of quantum information, the qubit, are characterized by two main parameters, that are fundamental indicators of their performance: the relaxation time $T_{1}$ and the coherence time $T_{2}^{\ast}$ [1, 2, 3, 4]. Both of these parameters must be optimized, and, in particular, decoherence, the process through which freely evolving qubits gradually lose stored information, represents one of the main limitations. Error mitigation techniques, in particular quantum error correction (QEC) algorithms, have been developed, but their successful implementation requires that qubits have coherence times long enough for the QEC routines to be executed [5, 6, 7]. Sources of decoherence are random physical processes in the environment surrounding the quantum processor, but also noise and fluctuations in the signals used to control qubits. Therefore, specifically understanding and reducing the effects of noise affecting control signals on decoherence are fundamental challenges within the overall development of quantum computers. Here, we focus on the investigation of the action of such a noise, and, in particular, of phase noise, which is the prevalent nonideality (other forms of noise, such as thermal and shot noise, have in general a minor effect in this specific application). We will mainly deal with superconducting qubits, which represent one of the most promising implementations for quantum computers, due to their relatively long coherence times, as well as to the ease of manipulation [8, 9, 10, 11].
@@
Over the years, the need to control quantum states with high precision has led to the development of Quantum Control Theory [12, 13, 14, 15, 16, 17]. An important aspect of this theory is the use of spectral overlap techniques to analyze and reduce the effect of environmental noise on control protocols [18, 19]. This approach considers the spectral characteristics of both the noise, through its power spectral density, and of the control pulses or pulse sequences, through the associated filter functions in the Fourier domain. Based on these assumptions, the theory of Dynamical Error Suppression (DES) was developed, which exploits a sequence of appropriately designed pulses to mitigate the effects of dephasing [20, 21, 22, 23, 24, 25, 26].
@@
A few years ago, the approach based on filter transfer functions was extended to the analysis of the impact, on the coherence and operational fidelity of qubits, of the phase fluctuations (phase noise) associated to the reference oscillators for control pulse generation. In particular, in Ref. [27] H. Ball et al. extended the usage of the filter function formalism to the study of the effect of the phase fluctuations of the reference clock by shifting such fluctuations onto the qubit. The approach developed by Ball et al. was later used in [28] to study phase noise and wide-band additive noise of a realistic oscillator, and, more recently, by Matsuoka et al. [29] to investigate the effect on fidelity of the phase noise due to the cryo CMOS oscillator. Indeed, the choice of the reference frame is arbitrary, and, therefore, one can assume the local oscillator frame as the reference; thus, phase fluctuations of the local oscillator will appear as opposite phase oscillations of the qubit.
@@
Our approach is based on a two-step method: first, we generate a pseudorandom sequence of phase values distributed in agreement with the assumed (or measured) noise Power Spectral Density (PSD); then, such a sequence is used as an input for a numerical simulation of the evolution in time of the qubit state, performed with Qiskit-Dynamics [30]. The final state reached by the qubit can then be compared with the one from an ideal, noiseless evolution. In the literature, alternatives to Qiskit-Dynamics exist to model the qubit evolution, such as QuTip [31], which, for example, was used by Tong et al. [32] to study the effect of constant phase and frequency shifts and of other errors in the control signal. Simulations are repeated for different realizations of the same phase noise process (changing the seed used to generate the pseudorandom sequence) and averaged to obtain an estimate for the fidelity achievable with each considered choice of noise power spectral density.
Method
The corresponding finite-length FIR impulse response is then:
$h[n]={}$
In our numerical procedure, we define the amplitude function $A_{d}(\omega)$ as the square root of the desired noise PSD, set $\psi_{0}=0$, and we choose even values for the order $M$ (Type I filter). This numerical procedure, shown in more detail in Fig. 1, allows us to obtain a phase noise process with the desired PSD, just by modeling the shape of the FIR filter frequency response. To accelerate the computation of the convolution in the time domain between the white Gaussian process and the impulse response of the filter, the Fast Fourier Transform (FFT) is leveraged to convert the operation into a multiplication in the frequency domain, performing proper zero-padding to ensure that the vector lengths are powers of 2 [34]. Once the baseband phase noise values with specific amplitude and spectral characteristics are obtained, they are used directly as the phase $\phi_{j}$ at time $t_{j}$ of the carrier for the control pulses ($s(t_{j})=\Re\{f(t_{j})\exp[i(2\pi\nu t_{j}+\phi_{j})]\}$, where $f(t_{j})$ is the pulse envelope at $t_{j}$, and $\nu$ is the carrier frequency), in order to simulate their effect on qubit evolution. In all of our simulations we have considered a rectangular envelope for the control pulses.
@@ Acknowledgement
Data availability
The data are available from the authors upon request.
@@
Declarations
Ethics approval
Not applicable.
Funding
This work was partially supported by the Italian Ministry of University and Research (MUR) in the framework of the CrossLab and FoReLab projects (Departments of Excellence). Further support was provided by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS), under Contract No. 89243024CSC000002, and by QUART&T, a project funded by the Italian National Institute for Nuclear Physics (INFN) within the Technological and Interdisciplinary Research Commission (CSN5). Agata Barsotti was supported by a PhD scholarship funded by the European Union – NextGenerationEU under the National Recovery and Resilience Plan (NRRP), within the project “PNRR - HPC, Big Data and Quantum Computing (CN1 Spoke 10: Quantum Computing)”.